OpenStudy (anonymous):

Charlotte purchased a pool for $7,680 using a six-month deferred payment plan with an interest rate of 20.45%. She did not make any payments during the deferment period. What will Charlotte’s monthly payment be if she must pay off the pool within three years after the deferment period? $213.33 $236.10 $317.82 $371.83

OpenStudy (amistre64):

take the initial balance and compound it monthly for 6 months

OpenStudy (amistre64):

then we can use that balance to determine the Payment for a 3 year payoff

OpenStudy (anonymous):

7680(1+0.2045/6)^3 ?

OpenStudy (amistre64):

/12

OpenStudy (amistre64):

and ^6

OpenStudy (anonymous):

With that I got 8499.51

OpenStudy (amistre64):

7680(1+0.2045/12)^6 = balance to pay off

OpenStudy (amistre64):

now we take that balance and our 1+r/12 and run it for 3*12 periods of payments

OpenStudy (amistre64):

\[8500~k^{36}(\frac{1-k}{1-k^{36}})~;~k=1+\frac{.2045}{12}\]

OpenStudy (anonymous):

$236.10 ?

OpenStudy (amistre64):

i get more like 317.83

OpenStudy (anonymous):

Hm, maybe I left something out